OpenStudy (anonymous):

what is the oxidation number of Cl in Cl2O3

OpenStudy (anonymous):

No charge, so the oxidation numbers should add to 0. Oxygen is usually -2, so you have: \[(-2)\times 3 + (x) \times 2 = 0\] Or \[-6 +2x = 0\] And you just solve for x, which is the oxidation # of Cl.